OpenStudy (anonymous):

Difficult Integral! Ok maybe not, but could someone walk me through the steps of integrating something like cos(x^2) ? I know the derivative of something like this is found via chain rule... is there some analogue for integration that I can use here?

OpenStudy (amistre64):

just becasue we CAN integrate, does not mean that we can always integrate

OpenStudy (amistre64):

there are tricks that can be had for cos(x^2) that have to do with maths like the Gamma function and/or power series

OpenStudy (amistre64):

\[cos(u)=1-\frac{u^2}{2!}+\frac{u^4}{4!}-\frac{u^6}{6!}+...\]

OpenStudy (anonymous):

Yea, unless you're a college senior level math major or a grad student, you won't be asked to integrate things like cos(x^2).

OpenStudy (amistre64):

\[\int\sum_{n=0}^{inf}\frac{1}{(2n)!}x^{4n}dx\] might work out like this :) \[\sum_{n=1}^{inf}\frac{1}{(2n)!~(4n+1)}x^{(4n+1)}\]